OpenStudy (anonymous):

The sum of two numbers is 170 and their difference is 26. Find the two numbers.

OpenStudy (turingtest):

x+y=170 x-y=26 add eqns together: 2x=196 x=98 y=170-x=170-98=72

OpenStudy (anonymous):

98 and 72

OpenStudy (anonymous):

Medal please

OpenStudy (turingtest):

you should show your work to earn the medal though @WasimahP